OpenStudy (anonymous):

$6300 is invested, part of it at 11% and part of it at 7%. For a certain year, the total yield is $565.00. How much was invested at each rate?

OpenStudy (confluxepic):

Welcome to Open Study.

OpenStudy (confluxepic):

Try to find the the interest formula.

OpenStudy (confluxepic):

Exactly.

OpenStudy (anonymous):

But I already have the principle, I'm looking for the amount at each rate

OpenStudy (kl0723):

@aum @zepdrix @satellite73

OpenStudy (aum):

Assume "x" dollars is invested at 11%. Then (6300-x) dollars is invested at 7% Interest = P*r*t where P is the Prinicipal r is the rate of interest in decimal t is the number of years invested. Interest for "x" dollars invested at 11% for 1 year is: x * 0.11 * 1 = 0.11x Interest for "(6300-x)" dollars invested at 7% for 1 year is: (6300-x) * 0.07 * 1 = 441 - 0.07x Total interest = 0.11x + 441 - 0.07x = 565 solve for x.

OpenStudy (anonymous):

thanks for all your work but it's not looking for interest. I really appreciate it.

OpenStudy (aum):

Once you solve for x, that is the amount invested at the 11% rate. (6300-x) is the amount invested at the 7% rate.

OpenStudy (aum):

Total interest = 0.11x + 441 - 0.07x = 565 (0.11x - 0.07x) = 565 - 441 0.04x = 124 x = 124 / 0.04 = 3100 6300-x = 6300-3100 = 3200 Therefore, $3,100 was invested at the 11% rate and $3,200 was invested at the 7% rate.